OpenStudy (anonymous):

Need help in solving this question 2 sqrt x+5 = x-3

OpenStudy (anonymous):

\[2\sqrt{x+5}=x-3\]

OpenStudy (anonymous):

on squaring u will get a quadratic equation \[x ^{2}-10x-11\]

OpenStudy (anonymous):

now solve this to get values of x and substitute these values on both sides the value which gives the same answer on both sides is ur value

OpenStudy (anonymous):

I got x =11 as the only solution when substituting is that correct?

OpenStudy (anonymous):

yes as x=-1 will give a complex no

OpenStudy (anonymous):

OK sweet, thanks for your help again

OpenStudy (anonymous):

ur welcome